# ProReason: Multi-Modal Proactive Reasoning with Decoupled Eyesight and Wisdom

### v2

@@ 6 Limitations
Considering that multiple agents are engaged in the answering process, we try to find instances where contradictory information is provided by different agents, especially the vision expert and reasoning expert. However, as shown in Figure 6 and 7, when one agent (e.g., the vision expert) makes an error and the referee even hints at a possible mistake, other agents (e.g., the reasoning expert) tend to adhere to the available information instead of questioning it. This tendency results in a failure to find cases with contradictory information, and also highlights the importance of a reflection mechanism (Ji et al., 2023) in agent collaboration, which is left for future exploration.
@@ A.1 Related work
Recently, large vision-language models (LVLMs) (Bai et al., 2023; Chen et al., 2023; Liu et al., 2024b) have garnered widespread attention and demonstrated remarkable advancements in understanding and generating multi-modal contents. In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c, b, 2024a; Li et al., 2024a, b) and InternVL (Chen et al., 2024b) families, have been extensively developed. In the closed-source domain, proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini Pro 1.5 (Reid et al., 2024) have also achieved significantly success. Despite these advancements, existing LVLMs still encounter challenges in effectively integrating visual understanding with textual reasoning capabilities simultaneously. This limitation is particularly evident in their diminished attention to image content during visual reasoning process, such as chart interpretation and visual math reasoning, leading to degraded performance (Ghosh et al., 2024) and motivating more effective solutions.
Passive Visual Reasoning.
Extracting information from images into text can effectively assist LVLM in performing visual reasoning tasks. Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024) first describes the image before prefixing this description to the prompt, assisting LVLMs on visual reasoning tasks. Furthermore, Compositional Chain-of-Thought (CCoT) (Mitra et al., 2024) directs LVLMs to generate scene graphs (SGs) that serve as a bridge between the visual and textual domains, aiding LVLMs in subsequent tasks. However, most of these methods employ a question-agnostic and reasoning-free visual extraction process, where image descriptions are not tailored to specific questions, and no reasoning is applied to infer additional information for improved descriptions. These “passive” approaches lead to the inclusion of irrelevant or redundant information, ultimately degrading performance. In contrast, ProReason adopts question-oriented agents to collect necessary and sufficient information, effectively circumventing these drawbacks.
@@
Multi-step reasoning frameworks have been developed for LLMs to achieve better performance by breaking down complex questions into easier ones (Pan et al., 2024). As a representative method, Chain-of-Thought (CoT) (Wei et al., 2022) enhances the arithmetic and commonsense reasoning capabilities by explicitly generating intermediate reasoning steps before concluding the final answers. Tree-of-Thoughts (ToT) (Yao et al., 2024) further refines the CoT mechanism by allowing LLMs to consider multiple reasoning paths and do self-assessment before making decisions. Considering that the inherent knowledge of LLMs may not be sufficient to complete tasks, ReAct (Yao et al., 2022) integrates information retrieval into the reasoning chains, enabling models to pause to verify results and determine whether additional information is needed before proceeding. Nevertheless, multi-step reasoning frameworks designed for text-only LLMs are not fully applicable to visual reasoning tasks, and may even impair the performance of LVLMs (Ghosh et al., 2024).
@@ A.2 Dataset
To thoroughly assess the performance of our framework, we have carried out experimental evaluations using four benchmark datasets: the Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a), the cross-disciplinary Massive Multi-modal Understanding and Reasoning benchmark (MMMU) (Yue et al., 2023), the visual mathematical reasoning assessment MathVista (Wang et al., 2024), and the multimodal illusion detection benchmark HallusionBench (Liu et al., 2023a). Each of these benchmarks necessitates strong visual reasoning capabilities for successful task completion, and we provide concise descriptions below:
@@
A.3 Supplementary Results and Analysis
A.3.1 Which One is More Crucial: Visual Understanding or Textual Reasoning?
@@
Textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, although both are important. As shown in Table 3, replacing either the Vision Expert or the other agents with the more capable GPT-4o-mini achieves consistent performance enhancement, highlighting the significance of both capabilities. However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert. This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Sec. 4.3 that identifies the Summarizer and Referee as the most crucial sub-agents.
@@
A.3.2 Referee’s Dispel of Hallucinations
Dataset Attempt Allowances
1 3 5
MMMU 60.1 59.9 61.6
HallusionBench 57.1 58.9 59.9
A.3.3 Performance of ProReason on VQA task
ProReason also excels in conventional VQA tasks. We perform an evaluation on the VQA benchmark A-OKVQA dataset (Schwenk et al., 2022). The visual complexity of A-OKVQA is comparable to MME, but its reasoning requirements are less demanding. Table 5 illustrates the performance of ProReason, powered by GPT-4o-mini, on the A-OKVQA, MME, and MMMU benchmarks. ProReason achieves the highest performance across all three benchmarks, indicating the positive impact of ProReason on any form of reasoning-involved multi-modal tasks.
ProReason is particularly effective on datasets with greater complexity due to its decoupled capabilities and LLM-assisted reasoning. Table 5 demonstrates that ProReason outperforms the "Direct" method by 2.7%, 11.2%, and 13.2% on A-OKVQA, MME, and MMMU, respectively, as the visual and reasoning complexity increases. This underscores the advantage of ProReason in handling more challenging tasks, making it a robust choice for complex problem-solving scenarios.
Table 5: Performance of multiple approaches with GPT-4o-mini across three visual benchmarks.
GPT-4o-mini Dataset
A-OKVQA MME MMMU
Direct 78.9 79.2 48.4
VDGD 79.4 82.3 51.4
CCoT 79.2 80.8 54.2
COT 80.9 87.8 58.5
ReAct 80.6 87.3 54.8
ProReason 81.3 (+2.7) 90.4 (+11.2) 61.6 (+13.2)
A.3.4 Evaluation of ProReason based on LLMs
In Sec. 2.1, we utilize a pipeline based on LLMs to evaluate the correlation between model responses and image content. We aim to apply this effective method once more to assess the performance of ProReason. Initially, we adopted the reasoning chains generated by GPT-4o-mini (CoT) as the standard answers, given its superior performance with a score of 58.5 on the MMMU benchmark, compared to other models. However, due to ProReason’s enhancement of GPT-4o-mini’s score to 61.6, the use of GPT-4o-mini’s answers became inadequate. Consequently, we opted for GPT-4o to generate the standard answers, achieving a score of 69.1 on MMMU.
As shown in Table 6, we evaluate the reasoning process of GPT-4o-mini (CoT), the Memory of ProReason (ProReason-Memory), and the reasoning process of Summarizer (ProReason-Summarizer), which generates the final answer of ProReason with the prompt template in Figure 12. This assessment focuses on three key metrics: the relevance to standard answers (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI↓), where arrows indicate the directions of improvement.
GPT-4o-mini Metrics
RE ↑ RI ↓ MI ↓
COT 4.67 3.33 1.40
ProReason-Memory 4.83 3.66 1.17
ProReason-Summarizer 4.83 2.88 1.33
A.3.5 Frequency of selection of various experts
As listed in Table 7, we evaluate the frequency of the Dispatcher choosing the Vision Expert or Reasoning Expert on both MME and MMMU benchmarks, with MMMU requiring higher visual and reasoning abilities. Specifically, compared to MME, the frequencies for both the Vision and Reasoning Experts are higher on the MMMU benchmark, aligning with their difficulty levels. Together with the results in Table 1 of our submission, ProReason can adaptatively increase the frequencies of experts, and provide consistent performance improvements (i.e., 11.2% and 13.2%). Despite the lower frequency of the reasoning expert, the significant performance enhancement highlights the importance of LLM-assisted reasoning capabilities for reasoning-essential questions. Additionally, the frequency of the Vision expert exceeding 1 underscores the importance of referees, which controls the loop to call experts multiple times, alleviating the issue of insufficient information.
Table 7: Frequency of the Dispatcher choosing the Vision Expert or Reasoning Expert on both MME and MMMU benchmarks.
Dataset GPT-4o-mini
Vision Expert Reasoning Expert
MME 1.16 0.12
MMMU 1.64 0.38
Model Method
Direct CoT VDGD
Llama3-LLaVA-NeXT-8B 41.8 41.5 42.3
LLaVA-OneVision-Qwen2-7B-OV 49.1 46.5 49.7
Qwen-VL-Chat 36.2 35.4 37.1
Table 9: Image Relevance Score of Chain-of-Thought reasoning traces for “True” and “False” responses of three LVLMs, respectively.
Model Image Relevance Score
True False
Llama3-LLaVA-NeXT-8B 3.70 3.36
LLaVA-OneVision-Qwen2-7B-OV 4.27 3.67
Qwen-VL-Chat 3.63 2.80
Score Llama3-LLaVA-NeXT-8B
True False
Detail Level 4.43 3.93
Question Relevance 3.87 3.30
Reasoning Effective Info Inclusion 3.91 3.57
Table 11: Average token consumption of ProReason and ReAct with GPT-4o-mini model on the MME and MathVista benchmarks.
Method MME MathVista
Input Output Input Output
ProReason 1286.8 327.2 2238.6 788.6
ReACT 1645.0 197.0 3092.8 845.1
Model Agent Dataset
MME MMMU
GPT-4o-mini 90.4 61.6
Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
Vision Expert 84.7 (-5.7) 54.5 (-7.1)
Reasoning Expert 88.7 (-1.7) 60.2 (-1.4)
Referee 89.6 (-1.1) 51.5 (-10.1)
Summarizer 84.2 (-6.2) 51.0 (-10.6)
@@ A.3.5 Frequency of selection of various experts
Vision & Reasoning Expert,
@@
Table 14: Performance of multiple approaches with two base models across four visual reasoning benchmarks. Different from Table 1, MME scores are calculated with the official method here.
@@
MME MMMU MathVista HallusionBench
Llama3-LLaVA- NeXT-8B Direct 35.9 41.8 37.1 45.8
VDGD 37.4 42.3 36.1 44.2
CCoT 35.4 40.5 36.8 37.4
CoT 25.8 41.5 35.9 43.1
ReAct 38.5 46.7 31.7 43.6
ProReason 41.6 50.5 38.8 50.9
GPT-4o-mini Direct 55.2 48.4 53.0 56.0
VDGD 58.4 51.4 51.2 52.4
CCoT 53.5 54.2 53.6 56.7
CoT 72.9 58.5 53.8 56.3
ReAct 71.2 54.8 49.3 51.1
ProReason 77.9 61.6 54.9 59.9
Table 15: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks. Different from Table 2, MME scores are calculated with the official method here.
Vision Expert Textual Sub-Agents Method Dataset
MME MMMU MathVista Hallu.
Llama3- LLaVA- NeXT-8B Llama3-LLaVA- NeXT-8B ReAct 385 46.7 31.7 43.6
ProReason 416 50.5 38.8 50.9
Qwen2.5-72B- Instruct ReAct 428 50.4 34.6 40.4
ProReason 603 56.8 48.8 52.3
GPT-4o-mini ReAct 473 48.4 36.2 46.7
ProReason 613 54.5 41.7 53.1
Table 16: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks. Different from Table 3, MME scores are calculated with the official method here.
Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
Textual Sub-Agents Vision Expert All Sub-Agents COT Direct
MME 613 512 416 286 359
MMMU 54.5 53.4 52.5 41.5 41.8
Table 17: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks. For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini. Different from Table 13, MME scores are calculated with the official method here.
Model Agent Dataset
MME MMMU
GPT-4o-mini 779 61.6
Llama3- LLaVA- NeXT-8B Dispatcher 746 60.9
Vision Expert 613 54.5
Reasoning Expert 743 60.2
Referee 762 51.5
Summarizer 641 51.0
Table 18: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments. Different from Table 13, MME scores are calculated with the official method here.
GPT-4o-mini Dataset
MMMU MME
ProReason 61.6 779
CoT 58.5 729
Vision & Reasoning
Expert Integration
58.6 673
Dispatcher,
Vision & Reasoning Expert,
& Referee Integration
57.8 714
All Sub-Agents Integration 56.8 711
A.4 Demonstrative Examples

### v3

@@ 6 Limitations
Considering that multiple agents are engaged in the answering process, we try to find instances where contradictory information is provided by different agents, especially the vision expert and Insight Expert. However, as shown in Figure 6 and 7, when one agent (e.g., the vision expert) makes an error and the referee even hints at a possible mistake, other agents (e.g., the Insight Expert) tend to adhere to the available information instead of questioning it. This tendency results in a failure to find cases with contradictory information, and also highlights the importance of a reflection mechanism (Ji et al., 2023) in agent collaboration, which is left for future exploration.
We also examined existing multi-step reasoning frameworks,
such as ReAct (Yao et al., 2022), ToT (Yao et al., 2024), and Insight-V (Dong et al., 2024), aiming to find solutions for resolving accumulated errors and contradictory information. However, we found that these approaches may also fail to identify effective solutions and do not discuss the impact of the aforementioned drawbacks. We sincerely WELCOME any constructive discussions regarding accumulated errors and contradictory information! Besides, ProReason demonstrates significantly improved overall performance, suggesting fewer errors made by ProReason in the whole task distribution.
@@ A.1 Related work
Recently, large vision-language models (LVLMs) (Yang et al., 2024; Bai et al., 2023; Chen et al., 2023; Liu et al., 2024b) have garnered widespread attention and demonstrated remarkable advancements in understanding and generating multi-modal contents. In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c, b, 2024a; Li et al., 2024a, b) and InternVL (Chen et al., 2024c) families, have been extensively developed. In the closed-source domain, proprietary models such as GPT-4o [^5] https://openai.com/index/hello-gpt-4o/ and Gemini Pro 2.5666https://deepmind.google/technologies/gemini/pro/ have also achieved significantly success. Additionally, multi-agent frameworks like VipAct (Zhang et al., 2024c) have been developed to improve LVLMs’ perception of visual details. Despite these advancements, existing LVLMs still encounter challenges in effectively integrating visual understanding with textual reasoning capabilities simultaneously. This limitation is particularly evident in their diminished attention to image content during visual reasoning process, such as chart interpretation and visual math reasoning, leading to degraded performance (Liu et al., 2023a; Ghosh et al., 2024) and motivating more effective solutions.
@@
Multi-step reasoning frameworks improve LLM performance by breaking down complex tasks. Chain-of-Thought (CoT) (Wei et al., 2022) enhances reasoning via explicit intermediate steps, demonstrating effectiveness in both textual and visual tasks (Zhang et al., 2024b), while Tree-of-Thoughts (ToT) (Yao et al., 2024) extends this by evaluating multiple reasoning paths. ReAct (Yao et al., 2022) integrates dynamic knowledge retrieval during reasoning. In visual reasoning, several frameworks assist LVLM by extracting image information into text. Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024) describes the image and appends this description to the prompt, aiding LVLMs in visual reasoning tasks. Compositional Chain-of-Thought (CCoT) (Mitra et al., 2024) guides LVLMs to create scene graphs (SGs) that link visual and textual domains, supporting subsequent tasks. However, these methods often use a question-agnostic, reasoning-free visual extraction process, resulting in irrelevant or redundant information. Insight-V (Dong et al., 2024) trains LVLM with a multi-agent system, but all agents rely on a single LVLM, blending visual understanding with textual reasoning and failing to address the challenge of effectively managing both. Unfortunately, this method does not release its code and prompts, resulting in low reproducibility and making comparison difficult. In response to these drawbacks, we introduce ProReason, which decouples visual reasoning tasks into proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and makes all prompts available. By leveraging the strengths of the decoupled system, ProReason effectively integrates existing powerful LLMs to achieve high-performance visual reasoning and successfully transfers this capability to downstream task models.
@@ A.2 Dataset
To thoroughly assess the performance of our framework, we have carried out experimental evaluations using four benchmark datasets: the Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023), the cross-disciplinary Massive Multi-modal Understanding and Reasoning benchmark (MMMU) (Yue et al., 2023), the visual mathematical reasoning assessment MathVista (Wang et al., 2024), and the multimodal illusion detection benchmark HallusionBench (Liu et al., 2023a). Each of these benchmarks necessitates strong visual reasoning capabilities for successful task completion, and we provide concise descriptions below:
@@
A.3 Baselines
  3. ReAct. ReAct is an LLM-specific agent framework, which performs tasks by alternating between reasoning and execution behaviors. To extend it to multi-modal domain, we use two LVLMs to perform both steps, and rename them as the Vision and Insight Experts, respectively. This aligns with our notions for easy understanding, and is shown in Figure 1.
  4. VDGD. VDGD involves two main steps: initially, LVLMs generate detailed image captions, which are then incorporated into prompts to aid inference. During the inference process, VDGD also utilizes a formula based on Kullback-Leibler divergence to select tokens that minimally deviate from the description, thereby enhancing the relevance of the model’s reasoning to the image [^7] Since we cannot obtain the tokens output by GPT-4o-mini, we omit the step of selecting the token with the smallest deviation from the image description when implementing VDGD for GPT-4o-mini..
  6. R1-Onevision. R1-Onevision-7B is an advanced multimodal reasoning model based on Qwen2.5-VL-7B-Instruct that has deep thinking capabilities akin to Deepseek-R1 by transforming images into structured textual representations and employing a training framework that merges supervised fine-tuning with reinforcement learning.
A.4 Model Training
[…]
A.5 Supplementary Results and Analysis
Model Method
Direct CoT VDGD
Llama3-LLaVA-NeXT-8B 41.8 41.5 42.7
Qwen2.5-VL-7B-Instruct 51.8 52.2 52.6
Score Llama3-LLaVA-NeXT-8B
True False
Detail Level 4.43 3.93
Question Relevance 3.87 3.30
Reasoning Effective Info Inclusion 3.91 3.57
A.5.1 Relative Importance of Sub-agents
To assess the importance of each sub-agent within the ProReason framework for visual reasoning tasks, we design five scenarios where Llama3-LLaVA-NeXT-8B acts as Dispatcher, Vision Expert, Insight Expert, Referee, or Summarizer, respectively, while the other sub-agents are powered by GPT-4o-mini. Given that Llama3-LaVA-NeXT-8B exhibits weaker visual understanding and textual reasoning capabilities than GPT-4o-mini, the more significant the performance drop incurred by replacing a sub-agent with Llama3-LaVA-NeXT-8B is, the more important that sub-agent is. Here we primarily consider the MME and MMMU benchmarks due to their comprehensive question coverage. The experimental results are presented in Table 6.
Relatively, Dispatcher and Insight Expert are the least essential sub-agents. Specifically, despite a decline, these two sub-agents exhibit significantly less performance degradation than other sub-agents. This can be attributed to the easier task of the Dispatcher, which requires minimal textual reasoning capabilities, and the infrequent calls of the Insight Expert, which is only activated when additional information needs to be inferred—a situation that is rare in current benchmarks. Besides, both sub-agents operate within the acquisition loop, allowing for greater error tolerance. Even if some error occurs, subsequent iterations can compensate for the missing information.
In summary, each sub-agent contributes to the performance of ProReason, underscoring their necessity. Relatively, the Summarizer and Referee are the most critical sub-agents, while the Dispatcher and Insight Expert have the least impact.
Model Agent Dataset
MME MMMU
GPT-4o-mini 90.4 61.6
Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
Vision Expert 84.7 (-5.7) 54.5 (-7.1)
Insight Expert 88.7 (-1.7) 60.2 (-1.4)
Referee 89.6 (-1.1) 51.5 (-10.1)
Summarizer 84.2 (-6.2) 51.0 (-10.6)
A.5.2 Which One is More Crucial: Visual Understanding or Textual Reasoning?
@@
Textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, although both are important. As shown in Table 7, replacing either the Vision Expert or the other agents with the more capable GPT-4o-mini achieves consistent performance enhancement, highlighting the significance of both capabilities. However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert. This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Section A.5.1 that identifies the Summarizer and Referee as the most crucial sub-agents.
@@
A.5.3 The Critical Implication of Decoupling
  1. Merge the Vision Expert and Insight Expert into a single sub-agent to examine the necessity of modality decoupling during sub-task execution.
  2. Integrate the Dispatcher, Vision Expert, Insight Expert, and Referee as a unified agent to verify the essentiality of the process design for the Proactive Visual Perception stage.
Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks. As shown in Table 8, merging Vision and Insight Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs. CoT’s 58.5). Although agent merging also causes performance declines on MME, the merged versions still outperform CoT. Notably, given MMMU’s substantially higher complexity compared to MME, these findings reveal that the decoupling of visual-textual processing fundamentally drives its performance gains in complex scenarios.
@@ A.5.3 The Critical Implication of Decoupling
Vision & Insight Expert,
@@
A.5.4 Reasoning Process Evaluation of ProReason
Inspired by Liu et al. (2023c), we design a pipeline using LLMs to further analyze the performance of ProReason. Given that GPT-4o achieved a 70.3% score on MMMU, significantly surpassing GPT-4o-mini’s 58.5% and ProReason driven by GPT-4o-mini’s 61.6%, we adopt GPT-4o’s CoT-based answers on MMMU as the standard answer. Using this standard answer, we evaluate the reasoning process of GPT-4o-mini (CoT), the Memory of ProReason powered by GPT-4o-mini (ProReason-Memory), and the reasoning process of ProReason’s Summarizer (ProReason-Summarizer) on MMMU. This assessment focuses on three key metrics: the relevance to standard answer (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI↓), where arrows indicate the directions of improvement. The evaluation process is driven by GPT-4o and the prompt is shown in Figure 12.
GPT-4o-mini Metrics
RE ↑ RI ↓ MI ↓
COT 4.67 3.33 1.40
ProReason-Memory 4.83 3.66 1.17
ProReason-Summarizer 4.83 2.88 1.33
A.5.5 Referee’s Dispel of Hallucinations
Dataset Attempt Allowances
1 3 5
MMMU 60.1 59.9 61.6
HallusionBench 57.1 58.9 59.9
A.5.6 Frequency of selection of various experts
As listed in Table 11, we evaluate the frequency of the Dispatcher choosing the Vision Expert or Insight Expert on both MME and MMMU benchmarks, with MMMU requiring higher visual and reasoning abilities. Specifically, compared to MME, the frequencies for both the Vision and Insight Experts are higher on the MMMU benchmark, aligning with their difficulty levels. Together with the results in Table 1 of our submission, ProReason can adaptatively increase the frequencies of experts, and provide consistent performance improvements (i.e., 11.2% and 13.2%). Despite the lower frequency of the Insight Expert, the significant performance enhancement highlights the importance of LLM-assisted reasoning capabilities for reasoning-essential questions. Additionally, the frequency of the Vision expert exceeding 1 underscores the importance of referees, which controls the loop to call experts multiple times, alleviating the issue of insufficient information.
Table 11: Frequency of the Dispatcher choosing the Vision Expert or Insight Expert on both MME and MMMU benchmarks.
Dataset GPT-4o-mini
Vision Expert Insight Expert
MME 1.16 0.12
MMMU 1.64 0.38
A.5.7 Performance Evaluation of ProReason on Supplementary Datasets
To further assess the effectiveness of ProReason across diverse visual tasks, we expand upon the four datasets introduced in Section 4.1 by incorporating three additional benchmarks: MathVerse (Zhang et al., 2024a), MMStar (Chen et al., 2024a), and A-OKVQA (Schwenk et al., 2022). MathVerse focuses on visual reasoning within mathematical contexts, while MMStar highlights visual dependency and data reliability. A-OKVQA, in contrast, serves as a knowledge-based VQA dataset requiring minimal reasoning. For MathVerse, we utilize the Text Lite subset, which prioritizes visual reasoning by minimizing textual content. In MMStar, subsets related to coarse and fine-grained perception that are unrelated to reasoning are excluded. Table 13 presents ProReason’s performance on each dataset, showcasing significant improvements across various forms of multi-modal tasks and highlighting the effectiveness and generalizability of the perception and reasoning decoupling approach.
Table 12: Number of iterations required for each problem during the Proactive Visual Perception phase on MME and MathVista benchmarks.
Model Dataset
MME MathVista
GPT-4o-mini 1.28 2.13
Table 13: Performance of multiple approaches with GPT-4o-mini across 7 visual benchmarks. “Hallu.” is the abbreviation of HallusionBench. Based on the performance of the direct method, red and blue signify the improvement and degradation, respectively.
@@ A.5.7 Performance Evaluation of ProReason on Supplementary Datasets
MME MMMU MathVista Hallu. MathVerse MMStar A-OKVQA
GPT-4o-mini Direct 79.2 48.4 53.0 56.0 28.2 46.2 78.6
VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6) 30.1 (+1.9) 47.0 (+0.8) 79.4 (+0.8)
CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7) 29.2 (+1.0) 45.4 (-0.8) 79.2 (+0.6)
CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3) 28.9 (+0.7) 47.2 (+1.0) 80.9 (+2.3)
ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9) 30.4 (+2.2) 46.7 (+0.5) 80.6 (+2.0)
ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9) 31.6 (+3.4) 49.1 (+2.9) 81.3 (+2.7)
Average 84.63 54.82 52.63 55.4 29.7 46.9 80.0
Table 14: Different model configurations of Textual Sub-Agents for implementing LLM-assisted visual reasoning in ProReason. The abbreviations 4o-mini, Qwen72B, and Qwen3 refer to GPT-4o-mini, Qwen2.5-72B-Instruct, and Qwen3-32B, respectively. "Assisted" stands for LLM-assisted reasoning.
Model Method Vision Expert Summarizer Other Textual Sub-Agents
Llama3-LLaVA-NeXT-8B ProReason + 4o-mini Assisted Llama3-LLaVA-NeXT-8B GPT-4o-mini GPT-4o-mini
ProReason + Qwen72B Assisted Llama3-LLaVA-NeXT-8B Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct
Qwen2.5-VL-7B-Instruct ProReason + Qwen72B Assisted Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct
ProReason + Qwen3 Assisted Qwen2.5-72B-Instruct Qwen3-32B Qwen2.5-72B-Instruct
A.6 Demonstrative Examples